# Post-hoc Spurious Correlation Neutralization with Single-Weight Fictitious Class Unlearning

### v1

@@ Abstract
Neural network training tends to exploit the simplest features as shortcuts to greedily minimize training loss. However, some of these features might be spuriously correlated with the target labels, leading to incorrect predictions by the model. Several methods have been proposed to address this issue. Focusing on suppressing the spurious correlations with model training, they not only incur additional training cost, but also have limited practical utility as the model misbehavior due to spurious relations is usually discovered after its deployment. It is also often overlooked that spuriousness is a subjective notion. Hence, the precise questions that must be investigated are; to what degree a feature is spurious, and how we can proportionally distract the model’s attention from it for reliable prediction. To this end, we propose a method that enables post-hoc neutralization of spurious feature impact, controllable to an arbitrary degree. We conceptualize spurious features as fictitious sub-classes within the original classes, which can be eliminated by a class removal scheme. We then propose a unique precise class removal technique that employs a single-weight modification, which entails negligible performance compromise for the remaining classes. We perform extensive experiments, demonstrating that by editing just a single weight in a post-hoc manner, our method achieves highly competitive, or better performance against the state-of-the-art methods.
@@ 1 Introduction
Artificial Neural Networks (ANNs) that employ Empirical Risk Minimization (ERM) [37] are prone to correlating spurious features to target labels [12, 23, 41]. Residing in training data, such features often provide shortcuts to minimize loss, causing over-reliance of the model on them for inference [46, 43]. This leads to poor model generalization. Currently, the prevalent paradigm of suppressing learning of spurious feature-target label correlation - aka spurious correlation - is robust model learning [25, 27, 31, 19], which either requires a subsequent model retraining [22, 1], or training the model robustly right from scratch [9, 40].
Figure 1: Illustration of the adopted fictitious class perspective. Top: A non-robust classifier relies on incorrect set of high-level features - fictitious sub-classes in a class - to mis-associate male gender to non-blond hair. Bottom: Removing the undesired fictitious (sub-)class from the set enables robust classification.
@@
Another intriguing insight in the literature is presented by Eastwood et al. [10], who argue that spurious features are not entirely harmful. Aligned with [10], we posit that some apparent spurious correlations may even help model generalization, provided the right underlying data distribution. For instance, (spuriously) correlating a seagull with the background of sea might help correctly recognizing a bird over a sea as a seagull in a typical natural image setting. Clearly, spuriousness is a subjective notion, and techniques addressing spurious correlation should provide the ability to control the extent to which it can be neutralized for optimal model performance. Unfortunately, existing methods generally fail to explicitly account for such control.
In this work, we propose a post-hoc technique that enables neutralizing the contribution of a high-level feature to model prediction by an arbitrary degree. We conceptualize high-level features as fictitious sub-classes within the original class. Our method withdraws model attention from a selected fictitious sub-class to control the model behavior (see Fig. 1). Building on a strong theoretical foundation, we introduce a single-weight editing method to unlearn a fictitious class such that our post-hoc editing is applied to the network connection that contributes the most to the original model behavior for that class. To find that connection, we analyze class-specific activations and gradients of the model for the given class. In addition to being a unique post-hoc method to address spurious correlation, our technique also does not require group-labeled samples for editing, as often required by the existing methods [32, 9, 17].
@@
    We propose the first-of-its-kind post-hoc model unlearning technique to address spurious correlations. Our method edits only a single model weight to break-off prediction reliance on irrelevant high-level features.
@@
    With extensive experiments, we demonstrate state-of-the-art or comparable spurious correlation mitigation performance while being the only post-hoc method. Our method also does not require group information.
@@ 2 Related Work
Spurious Correlation Mitigation: To mitigate spurious correlation, early methods employed distributionally robust optimization (DRO) which utilizes group annotations to up-weight the worst-group loss during optimization [32, 15, 30, 44]. More recently, along similar lines, Deng et al. [9] proposed robust model learning with progressive data expansion. Though effective in terms of achieved worst-group performance, its strong dependence on group annotations limits the practicality of this approach. To mitigate the issue, other works have previously proposed using only limited amount of group annotated data [34, 27]. An extreme scenario is where there is no group annotation available at all. Methods aiming at such scenario [8, 45, 42] try to predict this information during the training process. Recently, Chakraborty et al. [4] utilized explainability heatmaps for clustering the groups. This solves the group information requirement problem, however; similar to all the methods mentioned above, [4] remains an ante-hoc technique.
@@
Finding Significant Connections: The seminal work of Optimal Brain Damage [20] motivated the exploration to rank the neural network connections based on their importance in the classification task. This helps in reducing memory footprint of the model by pruning the unimportant connections, which also leads to better generalization and faster inference. This research direction is still active, pursuing model efficiency and performance gains by identifying the subsets of most important network connections to retain [35, 39, 16, 33]. Our work is partially inspired by the counter-objective of seeking the most significant connection not to keep, to enable our unique type of unlearning.
@@ 3 Problem Definition
Consider a neural model representing a classification function $f_{\theta}(.):\mathcal{X}\rightarrow\mathcal{Y}$ with parameters $\theta\in\Theta$ to be trained on a dataset $\mathcal{D}_{tr}=\{(x_{i},y_{i})\}_{i=1}^{n}$, with training samples $x_{i}\in\mathcal{X}$ and their corresponding class labels $y_{i}\in\mathcal{Y}$. Let us denote a spurious feature by $a\in A$, where $A$ is the set of all presumed spurious features existing in $\mathcal{D}_{tr}$. For our problem, a group is defined using $a\in A$ and $y\in\mathcal{Y}$ as $g:=(a,y)\in A\times\mathcal{Y}$ s.t. $g\in\mathcal{G}$, where $\mathcal{G}$ is the set of all groups in $\mathcal{D}_{tr}$. To suppress spurious correlation, the commonly sought objective [27] is to minimize
$\mathcal{L}_{worst\_group}(\theta)=\max_{{}_{g\in\mathcal{G}}}\mathbb{E}_{(x,y,a)\sim P_{g}}[\ell(f_{\theta}(x),y)],$ (1)
where $\mathcal{L}_{worst\_group}$ is known as the worst group loss, $P_{g}$ is the group conditioned data distribution and $\ell(.)$ is the model prediction loss. Our specific objective can be further defined as follows
$\operatorname*{arg\,min}_{\theta^{*}\in\Theta}(\mathcal{L}_{worst\_group}(\theta))~{}\text{s.t.}~{}||\theta^{*}-\theta||_{0}\leq\delta.$ (2)
In Eq. (2), $\theta^{*}$ constitutes the sought vector of the model weights, ||.$||.||_{0}$ denotes $\ell_{0}$-pseudo norm that counts the non-zero elements of the vector, and $\delta\in\mathbb{Z}^{+}$ is a pre-defined positive integer. In this work, we focus on $\delta=1$ which enforces changing only a single model weight to suppress the spurious correlation impact on classification.
@@ 4 Methodology
Overview: Our method considers a high-level data feature as a sub-class within the actual class. Hypothetically, such features can be classified by another classifier as its targets, hence their treatment as a class is well-justified. However, since we do not aim to actually classify them, we see them as fictitious classes. This simple perspective allows us to treat spurious features as classes whose information can potentially be removed from the model by machine unlearning, thereby enabling neutralization of the undesired spurious correlations previously learned by the model.
For the fictitious class removal, we propose a unique post-hoc technique that aims at making the hyperplane involved in classifying a fictitious class orthogonal to its original state. Considering our objective in Eq. (2), this transformation needs precision to ensure minimal changes to the original model. To that end, we restrict our class removal to only a single weight, i.e., $\delta=1$, that is associated with the most significant connection in the neural network for the fictitious class. Such a connection should ideally be as exclusive as possible to the concerned class to minimize propagation of the editing effect to other classes. Hence, we also theoretically motivate and justify this exclusiveness for the connection identified in our approach.
A complete overview of the proposed method is given in Fig. 2
@@ 4.1 Association of Neural Connections to Classes
Here, we explore to what degree a connection in the neural network contributes to the classification process both generally and class-specifically. The forward and backward passes in the learning process are analyzed separately. We show that there are contrastive class associative properties in ANNs in the forward and backward passes.
As formally posited in Theorem 1 part (a), in a forward pass, the underlying graph edges that receive high activations from the samples of a small number of classes are more discriminative compared to those that get similar activations from a larger number of classes. Based on part (b) of Theorem 1, in backward pass, the edges with associated high magnitude gradients for samples from a fewer number of classes are less significant for classification as compared to those that have similar magnitude gradients from the samples of a larger number of classes. We provide a simplified illustration of the phenomenon in Fig. 3 where an extremal case for a three-class scenario is provided.
@@
Let $e_{ji}^{(l)}$ and $e_{qp}^{(l)}$ be two separate edges connecting nodes $n_{i}^{(l)}$ and $n_{p}^{(l)}$ in layer $l$ to nodes $n_{j}^{(l+1)}$ and $n_{q}^{(l+1)}$ in layer $l+1$ for a neural network being trained on dataset $\mathcal{D}_{tr}$, containing samples $\mathcal{S}=\{s_{1},\dots,s_{M}\}$ to classify the set into classes $\mathcal{C}=\{c_{1},\dots,c_{K}\}$.
@@
For conciseness, proofs of both parts of Theorem 1 are delegated to the supplementary material. Here, we only discuss an intuitive sketch of those proofs. For proving part (a), we first model the remaining uncertainty about a class given the activation of a neuron $n$ as conditional entropy $H(c|n)$. Then, we estimate the amount of information gain after observing the activation of the neurons $n_{i}^{(l)}$ and $n_{p}^{(l)}$. Our analytical expressions show that when $|\mathcal{C}_{1}|>|\mathcal{C}_{2}|$ the information gain for $n_{p}^{(l)}$ exceeds that of $n_{i}^{(l)}$. For part (b), we first define mutual discrimination (MD) for the concerned edges $e_{ji}^{(l)}$ and $e_{qp}^{(l)}$ MD is explained below. Analyzing the MD analytically, the constraint $|\mathcal{C}_{1}|>|\mathcal{C}_{2}|$ leads to the expression showing higher values for $e_{ji}^{(l)}$, which entails larger contribution of $e_{ji}^{(l)}$ to the classification task.
An edge receiving gradients from multiple classes must discriminate between them. Hence, considering its mutual discrimination (MD) ability for the classes is mandatory to analyze its contribution to decision making. To clarify the employed MD notion, in Fig. 4, we illustrate the quadratic relationship between the MD values and classes for which an edge might receive high magnitude gradients. We employ MD to prove Theorem 1 part (b).
The apparently contrastive association of neural connections to classes in forward and backward passes, as identified above, can be explained intuitively as follows. In the forward pass, high activations for a large number of classes means that the neuron is less helpful in discriminating between those classes. Hence, the associated connections do not contribute much to the classification. Conversely, high magnitude gradients in the backward pass emulate high sensitivity of the edges to the associated classes. It helps more in narrowing down the classification decisions when the sensitivity remains high for more classes.
@@
where $\frac{\partial{\mathcal{L}^{(s)}}}{\partial{e_{ji}^{(l)}}}$ is the gradient of loss function w.r.t. the edge $e_{ji}^{(l)}$ of the model receiving sample $s$ of class $c$ as the input. Strictly speaking, $\mathcal{G}_{ac.}^{c}$ is a scalar value here, we slightly abuse the term gradient for brevity.
@@
Specific Class Association Score: We are eventually also interested in finding which connections are more important in classifying certain classes. The CA-score in Eq. (5) provides a measure to associate a connection to the classes. To define specific class association (SCA) score for an edge, we use the CA-score to scale the product of the accumulative class-wise activations and gradients with $\Gamma_{CA}$.

### v2

# Single-weight Model Editing for Post-hoc Spurious Correlation Neutralization

@@ Abstract
[…]
Neural network training tends to exploit the simplest features as shortcuts to greedily minimize training loss. However, some of these features might be spuriously correlated with the target labels, leading to incorrect predictions by the model. Several methods have been proposed to address this issue. Focusing on suppressing the spurious correlations with model training, they not only incur additional training cost, but also have limited practical utility as the model misbehavior due to spurious relations is usually discovered after its deployment. It is also often overlooked that spuriousness is a subjective notion. Hence, the precise questions that must be investigated are; to what degree a feature is spurious, and how we can proportionally distract the model’s attention from it for reliable prediction. To this end, we propose a method that enables post-hoc neutralization of spurious feature impact, controllable to an arbitrary degree. We conceptualize spurious features as fictitious sub-classes within the original classes, which can be eliminated by a class removal scheme. We then propose a unique precise class removal technique that makes a single-weight modification, which entails negligible performance compromise for the remaining classes. We perform extensive experiments, demonstrating that by editing just a single weight in a post-hoc manner, our method achieves highly competitive, or better performance against the state-of-the-art methods. Our implementation will be made public after acceptance.
Index Terms:
model editing, machine unlearning, spurious correlation mitigation, post-hoc algorithms
@@ 1 Introduction
Artificial Neural Networks (ANNs) that employ Empirical Risk Minimization (ERM) [1] are prone to correlate spurious features to target labels [2, 3, 4, 5]. Residing in training data, such features often provide shortcuts to minimize loss, causing over-reliance of the model on them for inference [6, 7, 8, 9]. This leads to poor model generalization. Currently, the prevalent paradigm of suppressing learning of spurious feature-target label correlation - aka spurious correlation - is robust model learning [10, 11, 12, 13, 14], which either requires a subsequent model retraining [15, 16], or training the model robustly right from scratch [17, 18, 19, 20, 21].
Figure 1: Illustration of the adopted fictitious class perspective. Top: A non-robust Empirical Risk Minimization (ERM) classifier may rely on incorrect set of high-level features - fictitious sub-classes within a class - to mis-associate male gender to non-blond hair. Bottom: Removing the undesired fictitious (sub-)class from the set enables robust classification.
@@
Another intriguing insight in the literature is presented by Eastwood et al. [23], who argue that so-called spurious features are not entirely harmful. Aligned with [23], we posit that some apparent spurious correlations may even help model generalization, provided the right underlying data distribution. For instance, (spuriously) correlating a seagull with the background of sea might help correctly recognizing a bird over a sea as a seagull in a typical natural image setting. Clearly, spuriousness is a subjective notion, and techniques addressing spurious correlation should provide the ability to control the extent to which it can be neutralized for optimal model performance. Unfortunately, existing methods generally fail to explicitly account for such a control.
In this work, we propose a post-hoc technique that enables neutralizing the contribution of a high-level feature to model prediction by an arbitrary degree. We conceptualize high-level features as fictitious sub-classes within the original class. Our method withdraws model attention from a selected fictitious sub-class to control the model behavior (see Fig. 1). Building on a strong theoretical foundation, we introduce a single-weight editing method to unlearn a fictitious class such that our post-hoc editing is applied to the network connection that contributes the most to the original model behavior for that class. To find that connection, we analyze class-specific activations and gradients of the model for the given class. In addition to being a unique post-hoc method to address spurious correlation, our technique also does not require group-annotated samples for editing, as often required by some existing methods [24, 17, 22].
@@
    We propose a first-of-its-kind post-hoc model unlearning technique to address spurious correlations. Our method edits only a single model weight to break off prediction reliance on irrelevant high-level features in the input.
@@
    Our method operates without group label annotations and, through extensive experiments, demonstrates state-of-the-art or comparable performance in mitigating spurious correlations while being the only post-hoc method.
@@ 2 Related Work
Spurious Correlation Mitigation: To mitigate spurious correlation, early methods employed distributionally robust optimization (DRO), which utilizes group annotations to up-weight the worst-group loss during optimization [24, 25, 26, 27, 28]. More recently, along similar lines, Deng et al.  [17] proposed robust model learning with progressive data expansion. Though effective in terms of achieved worst-group performance, its strong dependence on group annotations limits its practicality. To mitigate the issue, other works proposed using only a limited amount of group annotated data [29, 11]. An extreme scenario is where there is no group annotation available at all. Methods aiming at such a scenario [30, 31, 32] try to predict this information during the training process. In [33] utilized explainability heatmaps for clustering the groups. This solves the group label annotations requirement problem, however; similar to all the methods mentioned above, [33] remains an ante-hoc technique. Recently, [34] aimed mitigation of spurious correlations for unbiased scene graph generation, and [35] explored the negative effect of foreground-background priors in open set recognition.
@@
Finding Significant Connections: The seminal work of Optimal Brain Damage [45] motivated the exploration to rank the neural network connections based on their importance in the classification task. This helps in reducing the memory footprint of the model by pruning the unimportant connections, which also leads to better generalization and faster inference. This research direction is still active, pursuing model efficiency and performance gains by identifying the subsets of most important network connections to retain [46, 47, 48, 49]. Our work is partially inspired by the counter-objective of seeking the most significant connection not to keep, to enable our unique type of unlearning.
@@ 3 Problem Definition
Consider a neural model representing a classification function $f_{\theta}(.):\mathcal{X}\rightarrow\mathcal{Y}$ with parameters $\theta\in\Theta$ to be trained on a dataset $\mathcal{D}_{tr}=\{(x_{i},y_{i})\}_{i=1}^{n}$, with training samples $x_{i}\in\mathcal{X}$ and their corresponding class labels $y_{i}\in\mathcal{Y}$. Let us denote a spurious feature by $\alpha\in A$, where $A$ is the set of all presumed spurious features existing in $\mathcal{D}_{tr}$. For our problem, a group is defined using $\alpha\in A$ and $y\in\mathcal{Y}$ as $g:=(\alpha,y)\in A\times\mathcal{Y}$ s.t. $g\in\mathcal{G}$, where $\mathcal{G}$ is the set of all groups in $\mathcal{D}_{tr}$. To suppress spurious correlation, the commonly sought objective [11] is to minimize
$\mathcal{L}_{worst\_group}(\theta)=\max_{{}_{g\in\mathcal{G}}}\mathbb{E}_{(x,y,\alpha)\sim P_{g}}[\ell(f_{\theta}(x),y)],$ (1)
where $\mathcal{L}_{worst\_group}$ is the worst group loss, $P_{g}$ is the group conditioned data distribution and $\ell(.)$ is the model prediction loss. Our specific objective is further defined as
[…]
In Eq. (2), $\theta^{*}$ constitutes the sought vector of the model weights, $||.||_{0}$ denotes $\ell_{0}$-pseudo norm that counts the non-zero elements of the vector, and $\delta\in\mathbb{Z}^{+}$ is a pre-defined positive integer. In this work, we focus on $\delta=1$ which enforces changing only a single model weight to suppress the spurious correlation impact on classification.
@@ 4 Methodology
Overview: Figure 2 shows an overview of our method which considers a high-level data feature as a sub-class within the actual class. Hypothetically, such features can be classified by another classifier as its targets, hence their treatment as a class is well-justified. However, since we do not aim to actually classify them, we see them as fictitious classes. This simple perspective allows us to treat spurious features as classes whose information can potentially be removed from the model by machine unlearning, thereby enabling neutralization of the undesired spurious correlations previously learned by the model.
For fictitious class removal, we propose a unique post-hoc technique that makes the hyperplane involved in classifying a fictitious class orthogonal to its original state. Considering our objective in Eq. (2), this transformation needs precision to ensure minimal changes to the original model. To that end, we restrict our class removal to modifying only a single weight, i.e., $\delta=1$, that is associated with the most significant connection in the neural network for the fictitious class. Such a connection should ideally be as exclusive as possible to the concerned class to minimize propagation of the editing effect to other classes. Hence, we also theoretically motivate and justify this exclusiveness for the connection identified in our approach.
@@ 4.1 Association of Neural Connections to Classes
Here, we explore to the degree to which a connection in the neural network contributes to the classification process both generally and class-specifically. The forward and backward passes are analyzed separately. We show that there are contrastive class associative properties in ANNs in the forward and backward passes.
As formally posited in Theorem 1 part (a), in a forward pass, the edges that receive high activations from the samples of a small number of classes are more discriminative compared to those that get similar activations from a larger number of classes. Based on part (b) of Theorem 1, in backward pass, the edges with associated high magnitude gradients for samples from a fewer number of classes are less significant for classification as compared to those that have similar magnitude gradients from the samples of a larger number of classes. We provide a simplified illustration of the phenomenon in Fig. 3 where an extremal case for a three-class scenario is provided.
@@
Let $e_{ji}^{(l)}$ and $e_{qp}^{(l)}$ be two separate edges connecting nodes $n_{i}^{(l)}$ and $n_{p}^{(l)}$ in layer $l$ to nodes $n_{j}^{(l+1)}$ and $n_{q}^{(l+1)}$ in layer $l+1$ for a neural network trained on dataset $\mathcal{D}_{tr}$, containing samples $\mathcal{S}=\{s_{1},\dots,s_{M}\}$ to classify the set into classes $\mathcal{C}=\{c_{1},\dots,c_{K}\}$.
@@
Proof.
(a) Let the conditional entropy $H(C|n)$ be the remaining uncertainty about the class given the activation of neuron $n$,
$H(\mathcal{C}|n)=-\sum_{c\in\mathcal{C}}p(c|n)\log{p(c|n)}.$ (3)
So, for the neurons $n_{i}^{(l)}$ and $n_{p}^{(l)}$ we have
$\begin{split}H(\mathcal{C}|n_{i}^{(l)})=-\sum_{c\in\mathcal{C}_{1}}p(c|n_{i}^{(l)})\log{p(c|n_{i}^{(l)})}\\ =-\sum_{|\mathcal{C}_{1}|}\frac{1}{|\mathcal{C}_{1}|}\log{\frac{1}{|\mathcal{C}_{1}|}}=\log|\mathcal{C}_{1}|\end{split}$ (4)
and
$\begin{split}H(\mathcal{C}|n_{p}^{(l)})=-\sum_{c\in\mathcal{C}_{2}}p(c|n_{p}^{(l)})\log{p(c|n_{p}^{(l)})}\\ =-\sum_{|\mathcal{C}_{2}|}\frac{1}{|\mathcal{C}_{2}|}\log{\frac{1}{|\mathcal{C}_{2}|}}=\log|\mathcal{C}_{2}|.\end{split}$ (5)
Next, we calculate the amount of information gain $\xi_{IG}$ after observing activation of the neurons $n_{i}^{(l)}$ and $n_{p}^{(l)}$
$\xi_{IG}(n_{i}^{(l)})=H(\mathcal{C})-H(\mathcal{C}|n_{i}^{(l)})=H(\mathcal{C})-\log|\mathcal{C}_{1}|,$ (6)
$\xi_{IG}(n_{p}^{(l)})=H(\mathcal{C})-H(\mathcal{C}|n_{p}^{(l)})=H(\mathcal{C})-\log|\mathcal{C}_{2}|.$ (7)
Since we have assumed in the theorem that $|\mathcal{C}_{1}|>|\mathcal{C}_{2}|$, from Eq. (6) and Eq. (7) we have
$\xi_{IG}(n_{i}^{(l)})<\xi_{IG}(n_{p}^{(l)}).$ (8)
So, receiving activations on $e_{ji}^{(l)}$ reveals less information for the model for discriminating between the classes.
(b) By definition, for a specific model parameter $e$, if the gradient of a class specific loss w.r.t. to that parameter $\frac{\partial{\mathcal{L}^{(c)}}}{\partial{e}}$ is high in magnitude, it shows that small changes in $e$ have significant impact on distinguishing between the class $c$ and other classes. If the parameter exhibits high gradients for $|\mathcal{C}_{k}|$ classes where $\mathcal{C}_{k}\subseteq\mathcal{C}$, we can calculate the mutual discrimination (MD) frequency ($\nu_{MD}$) based on the combination rule as $\binom{|\mathcal{C}_{k}|}{2}$. Computing this for edges $e_{ji}^{(l)}$ and $e_{qp}^{(l)}$, and simplifying, we get
$\nu_{MD}(e_{ji}^{(l)})=\frac{|\mathcal{C}_{1}|(|\mathcal{C}_{1}|-1)}{2},$ (9)
and
$\nu_{MD}(e_{qp}^{(l)})=\frac{|\mathcal{C}_{2}|(|\mathcal{C}_{2}|-1)}{2}.$ (10)
Based on the theorem assumption $|\mathcal{C}_{1}|>|\mathcal{C}_{2}|$, Eq. (9) and Eq. (10) give us
$\nu_{MD}(e_{ji}^{(l)})>\nu_{MD}(e_{qp}^{(l)}).$ (11)
So, as $e_{ji}^{(l)}$ takes role in more mutual discriminations, it contributes more to the classification task compared to $e_{qp}^{(l)}$. ∎
The apparently contrastive association of neural connections to classes in forward and backward passes, as identified above, can be explained intuitively as follows. In the forward pass, high activations for a large number of classes mean that the neuron is less helpful in discriminating between those classes. Hence, the associated connections do not contribute much to the classification. Conversely, high magnitude gradients in the backward pass emulate high sensitivity of the edges to the associated classes. It helps more in narrowing down the classification decisions when the sensitivity remains high for more classes (see Fig. 4).
Figure 5: Orthogonalization of the hyperplane classifying a class w.r.t. the most significant connection’s weight, i.e., $w_{1}$ - the only weight. It results in canceling the well-trained model’s discriminatory ability.
@@
where $\frac{\partial{\mathcal{L}^{(s)}}}{\partial{e_{ji}^{(l)}}}$ is the gradient of loss function w.r.t. the edge $e_{ji}^{(l)}$ of the model receiving sample $s$ of class $c$ as the input. More precisely, $\mathcal{G}_{ac.}^{c}$ is a scalar value here, and we generalized the term gradient for brevity.
@@
Specific Class Association Score: We are also eventually interested in finding which connections are more important in classifying certain classes. The CA-score in Eq. (14) provides a measure to associate a connection to the classes. To define a specific class association (SCA) score for an edge, we use the CA-score to scale the product of the accumulative class-wise activations and gradients with $\Gamma_{CA}$.